### v2

# Stability of optimal transport on metric measure spaces

Abstract
We prove a quantitative stability of Kantorovich potentials on non-smooth metric measure spaces with synthetic lower Ricci curvature bound, thereby confirming a recent conjecture of Kitagawa, Letrouit and Mérigot. Our proof, which employs the heat kernel-regularized $c$-transform, does not rely on linear structure or sectional curvature bounds, is new even in the smooth setting. As a corollary, we get a quantitative stability of optimal transport maps on Alexandrov spaces with lower curvature bound.
Keywords: optimal transport, heat kernel, Kantorovich potential, quantitative stability, metric measure space, curvature-dimension condition, Ricci curvature
MSC 2020: 53C23, 51F99, 49Q22
1 Introduction
1.1 Background and Motivation
Optimal transport, initiated by Monge [35] and reframed by Kantorovich [24], seeks the most efficient way to redistribute mass between two probability distributions. Precisely, given two probability measures $\rho$ and $\mu$ defined on Polish spaces $X$ and $Y$ respectively, and a cost function $c:X\times Y\rightarrow\mathbb{R}$, the Monge optimal transport problem aims to find a minimizer $T$, called an optimal transport map, of the following optimization problem among all measurable maps $H:X\to Y$ pushing $\rho$ forward to $\mu$:
$\int_{X}c(x,T(x))\,{\mathrm{d}}\rho(x)=\inf_{H_{\sharp}\rho=\mu}\int_{X}c(x,H(x))\,{\mathrm{d}}\rho(x).$ (MP)
The Monge problem is not always well-posed, since it prevents mass splitting. Its relaxation, called Kantorovich problem, optimizes over joint couplings instead of deterministic maps. By Kantorovich duality, the Kantorovich problem is equivalent to the following dual problem:
$\sup_{\phi(x)+\psi(y)\leq c(x,y)}\left\{\int_{X}\phi(x)\,{\mathrm{d}}\rho(x)+\int_{Y}\psi(y)\,{\mathrm{d}}\mu(y)\right\}.$ (KD)
The optimal functions $\phi,\psi$, which always exist under standard assumptions, are called Kantorovich potentials [2, 45].
In the Euclidean space, for $c(x,y)=\frac{1}{2}|x-y|^{2}$, Brenier’s landmark result [5] established that, for absolutely continuous source measures, the optimal transport map is unique and takes the form $T(x)=\nabla u(x)$ for a convex function $u$. This was later extended to Riemannian manifolds by McCann [32], who showed that for $c(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y)$, the optimal transport map is given by $T(x)=\exp_{x}(-\nabla\varphi(x))$ for a Kantorovich potential $\varphi$.
A fundamental question in both theoretical and applied contexts of optimal transport is the quantitative stability of optimal transport maps and Kantorovich potentials under perturbations of the target measure (see Letrouit’s lecture note [28] and the references therein). Based on recent breakthroughs in the quantitative stability of optimal transport on Euclidean spaces [11, 30], on boundaries of convex bodies [27, 20] and on Riemannian manifolds [26], Kitagawa–Letrouit–Mérigot [26, §1.2] conjecture that:
the quantitative stability results are also true in more general metric measure spaces with synthetic curvature bounds.
1.2 General Setting
We confirm the conjecture of Kitagawa–Letrouit–Mérigot in the following setting.
A. Metric measure spaces:
An Alexandrov space is a geodesic space of finite Hausdorff dimension and of curvature bounded from below (cf. [8]). An ${\rm RCD}(K,N)$ space is a metric measure spaces verifying the synthetic Riemannian curvature-dimension condition [43, 31, 3, 12]. An $n$-dimensional Alexandrov space with curvature bounded from below by $k$, equipped with its $n$-Hausdorff measure, is an ${\rm RCD}(k(n-1),n)$ space [41, 46]. After [5, 32], Gigli–Rajala–Sturm [15, Theorem 1.1] proved the existence and uniqueness of the optimal transport map on an $\rm{RCD}(K,N)$ space $(X,{\mathrm{d}},\mathfrak{m})$, for the quadratic cost $c(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y)$ and the source measure $\rho\ll\mathfrak{m}$.
B. Source measures:
In [30, Theorem 1.9], it has been shown that when the source measure $\rho$ is the uniform density on some non-John domain $S\subset\mathbb{R}^{n}$, then no quantitative stability estimates of the form
$\|\phi_{\mu}-\phi_{\nu}\|_{L^{2}(\rho)}\leq CW_{p}^{\alpha}(\mu,\nu).$
A bounded open subset $S$ of a metric space is called a John domain if there is a distinguished point $x_{0}\in S$ and a constant $\eta>0$ such that, for every $x\in S$, there is a rectifiable curve $\gamma\colon[0,\ell(\gamma)]\to S$ parametrized by arc length, such that $\gamma(0)=x$, $\gamma(\ell(\gamma))=x_{0}$, and
${\mathrm{d}}(\gamma(t),S^{c})\geq\eta t,~~~\forall t\in[0,\ell(\gamma)].$
John domains encompass many cases of interest, such as bounded Lipschitz domains, bounded domains satisfying a cone condition and certain fractal domains (see [7] for more discussions).
[…]
C. Kantorovich potentials:
It is also necessary to (zero-mean) normalize the Kantorovich potential $\phi$ from $\rho$ such that $\mathbb{E}_{\rho}(\phi)=\int_{S}\phi\,{\mathrm{d}}\rho=0$. Together with the uniqueness of the optimal transport map, such $\phi$ is unique and thus makes sense to talk about its stability.
1.3 Main Results
Our main theorem concerns the quantitative $L^{1}$ stability of Kantorovich potentials on ${\rm RCD}(K,N)$ spaces.
Theorem 1.1.
[…]
$\|\phi_{\mu}-\phi_{\nu}\|_{L^{1}(\rho)}\leq CW_{1}^{\frac{1}{2}}(\mu,\nu),$ (1.1)
where $\phi_{\mu}$ and $\phi_{\nu}$ are the Kantorovich potentials from $\rho$ to $\mu$ and $\rho$ to $\nu$ respectively.
In particular, if $(S,{\mathrm{d}},\mathfrak{m})$ is a compact ${\rm RCD}(K,N)$ space, $S$ is surely a John domain. So we have the following corollary.
Corollary 1.2.
Let $(X,{\mathrm{d}},\mathfrak{m})$ be a compact ${\rm RCD}(K,N)$ metric measure space. Let $\rho\in\mathcal{P}(X)$ be with $a_{1}\mathfrak{m}\leq\rho\leq a_{2}\mathfrak{m}$ for some positive constants $a_{1},a_{2}$. Then the conclusion of Theorem 1.1 holds.
Adapting the strategy of [26], we can also prove the stability of optimal transport maps on Alexandrov spaces.
Theorem 1.3.
Let $(X,{\mathrm{d}})$ be an $n$-dimensional Alexandrov space with curvature bounded from below by $k$, $\mathfrak{m}$ be the $n$-Hausdorff measure. Under the same assumption for $S,Y$ and $\rho$ as in Theorem 1.1, and if $S$ additionally has finite perimeter, then there exists a constant $C>0$, depending on $k,n,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),S$, such that for any $\mu,\nu\in\mathcal{P}(Y)$,
$\int_{S}{\mathrm{d}}^{2}(T_{\mu}(x),T_{\nu}(x))\,{\mathrm{d}}\rho(x)\leq CW_{1}^{1/6}(\mu,\nu),$ (1.2)
where $T_{\mu}$ and $T_{\nu}$ are the optimal transport maps from $\rho$ to $\mu$ and $\rho$ to $\nu$ respectively.
1.4 Strategy: heat kernel-regularized $c$-transform
Motivated by regularized $c$-transforms using Gibbs kernels $e^{-c(x,y)/\varepsilon}$ [26], entropic optimal transport [16, 17] and Varadhan’s formula
$\lim_{t\to 0}-t\log p_{t/2}(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y)=c(x,y),$
we make use of the following heat kernel-regularized $c$-transform  [^1] We are told by Luca Taminini that in solving the Schrödinger equation with the Cole–Hopf transform, a similar formula will occur.:
$\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\ni\psi\mapsto\Phi_{t}[\psi](x)=-t\log\int_{X}e^{\frac{\psi(y)}{t}}p_{t/2}(x,y)\,{\mathrm{d}}\mathfrak{m}(y).$
This approach allows us to bypass the low regularity of Kantorovich potentials in the non-smooth setting.
We define the heat kernel regularized Kantorovich functional as
$\mathrm{K}_{t}[\psi]:=\int_{S}\Phi_{t}[\psi]\,{\mathrm{d}}\rho.$
Similar to [26], the key in our proof is the strong concavity of the functional $\mathrm{K}_{t}$. To achieve this, we first derive a local strong concavity estimate using heat kernel estimate, then globalize the estimate on the support of the source measure using a Boman chain argument for John domains.
Unlike the regularized $c$-transform used in [26], the existence of the boundary of $Y$ may lead to the failure in our heat kernel regularization argument. This possibility is ruled out by using the measure concentration property of the heat kernel and by making a careful choice of Lipschitz extension.
Organization. This paper is structured as follows. In Section 2, we prove the quantitative stability of Kantorovich potentials on $\mathrm{RCD}(K,N)$ spaces. In Section 3, we establish the stability of optimal transport maps on Alexandrov spaces. The Appendix A contains technical lemmas about Poincaré inequalities.
Acknowledgement. The authors thank Nicola Gigli, Jun Kitagawa, Nan Li, Cyril Letrouit, Quentin Mérigot, Luca Tamanini for helpful discussions and suggestions on the bibliography.
2 Stability of Kantorovich potentials
2.1 Heat kernel estimate
We begin by recalling the short-time asymptotic behaviour of the heat kernel on $\rm RCD$} spaces. It was first studied by Varadhan on Riemannian manifolds [44], and is known as Varadhan formula today.
Lemma 2.1.
Let $p_{t}(x,y)$ be the heat kernel on an ${\rm RCD}(K,N)$ space $(X,{\mathrm{d}},\mathfrak{m})$. Then
$\lim_{t\rightarrow 0}-t\log p_{\frac{t}{2}}(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y),\quad\text{uniformly in}\quad x\in S,\,y\in Y.$ (2.1)
Proof.
By the heat kernel estimate [22, Theorem 1.2], for any $\epsilon>0$, it holds
$\frac{1}{C_{1}(\epsilon)\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)}$ $\exp\left(-\frac{{\mathrm{d}}^{2}(x,y)}{(4-\epsilon)t}-C_{2}(\epsilon)t\right)$ (2.2)
$\leq p_{t}(x,y)\leq$ $\frac{C_{1}(\epsilon)}{\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)}\exp\left(-\frac{{\mathrm{d}}^{2}(x,y)}{(4+\epsilon)t}+C_{2}(\epsilon)t\right).$
Then we have
$-t\log$ $\frac{C_{1}(\epsilon)}{\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)}+\frac{2{\mathrm{d}}^{2}(x,y)}{4+\epsilon}-\frac{C_{2}(\epsilon)}{2}t^{2}$ (2.3)
$\leq$ $-t\log p_{\frac{t}{2}}(x,y)\leq t\log C_{1}(\epsilon)\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)+\frac{2{\mathrm{d}}^{2}(x,y)}{4-\epsilon}+\frac{C_{2}(\epsilon)}{2}t^{2}.$
By Bishop–Gromov inequality [43, Theorem 2.3], we have
$\lim_{t\rightarrow 0}t\log\mathfrak{m}\left(B\big(y,\sqrt{\frac{t}{2}}\big)\right)=0,\quad\text{uniformly in }\,y\in Y.$ (2.4)
Letting $t\rightarrow 0$ in (2.3), we get the following uniform estimate:
[…]
Letting $\epsilon\rightarrow 0$, we prove the lemma. ∎
2.2 Heat kernel regularization
To establish the quantitative stability, we adopt a regularization technique similar to those in [36, 26] and [16, 17] (see also [9, 10, 34]).
For $\psi\in\mathop{\rm Lip}\nolimits(Y,{\mathrm{d}})$, define
$\bar{\psi}(y)=\sup_{z\in Y}\big\{\psi(z)-\text{Lip}(\psi)\,{\mathrm{d}}(z,y)\big\},\quad y\in X.$
[…]
For $t>0$, we define the heat kernel regularized $c$-transform as
$\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\ni\varphi\mapsto\Phi_{t}[\varphi](x):=-t\log\int_{X}e^{\frac{{\varphi}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y),$
and define the heat kernel regularized Kantorovich functional as
$\mathrm{K}_{t}[\varphi]:=\int_{S}\Phi_{t}[\varphi]\,{\mathrm{d}}\rho.$
For $\varphi\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$ and $x\in X$, we associate a probability measure
${\mathrm{d}}\mu_{x}^{t}[\varphi](y):=\frac{e^{\frac{\varphi(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\varphi(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}$
and define $\mu^{t}[\varphi]:=\int_{X}\mu_{x}^{t}[\varphi]\,{\mathrm{d}}\rho(x)$. This means, for any $v\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$, it holds
$\mathbb{E}_{\mu^{t}[\varphi]}(v)=\int_{X}\mathbb{E}_{\mu_{x}^{t}[\varphi]}(v)\,{\mathrm{d}}\rho(x).$ (2.6)
Next we prove the Sobolev regularity of $x\mapsto\mathbb{E}_{\mu_{x}^{t}[\varphi]}(v)$. We refer to [3] for Sobolev calculus on metric measure spaces.
Lemma 2.2.
For $t\in\big(0,\frac{D+1}{\sqrt{(N-1)|K|}}\wedge 1\big)$, $\psi\in\mathop{\rm Lip}\nolimits(Y,{\mathrm{d}})$ and $v\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$, there holds $\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$.
Proof.
Denote by $H_{t}f(x)=\int_{X}f(y)p_{t}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)$ the heat flow from $f$. Then
$\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)=\frac{\int_{X}v(y)e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}=\frac{H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})(x)}{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)}.$
Claim: $ve^{\frac{\psi_{*}}{t}},e^{\frac{\psi_{*}}{t}}\in L^{2}(\mathfrak{m})\cap L^{\infty}(\mathfrak{m})$.
Note that $|v(y)|\leq\sup_{Y}|v|+\mathop{\rm Lip}\nolimits(v)\mathop{\rm diam}\nolimits(Y)+\mathop{\rm Lip}\nolimits(v){\mathrm{d}}(y,Y)$ and $\psi_{*}(y)\leq\sup_{Y}\psi-(D+1){\mathrm{d}}(y,Y)$. So $ve^{\frac{\psi_{*}}{t}},e^{\frac{\psi_{*}}{t}}\in L^{\infty}(\mathfrak{m})$.
Fix $y_{0}\in Y$ and denote $B_{i}=\{y\in X:i\leq{\mathrm{d}}(y,Y)<i+1\},i\in\mathbb{N}$. It holds
$\int_{X}e^{\frac{2\psi_{*}}{t}}{\mathrm{d}}\mathfrak{m}$ $\leq\sum_{i=0}^{+\infty}e^{\frac{2\sup_{Y}\psi}{t}}e^{\frac{-2(D+1)i}{t}}\mathfrak{m}(B_{i})$ (2.7)
$\leq\sum_{i=0}^{+\infty}e^{\frac{2\sup_{Y}\psi}{t}}e^{\frac{-2(D+1)i}{t}}\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1)).$
Without loss of generality, we may assume $K<0$. By Bishop–Gromov inequality [43, Theorem 2.3], we have
$\mathfrak{m}\big(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1)\big)\leq c_{1}\int_{0}^{\mathop{\rm diam}\nolimits(Y)+i+1}\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s)\,{\mathrm{d}}s$
where $c_{1}=\frac{\mathfrak{m}(B(y_{0},\mathop{\rm diam}\nolimits(Y)))}{\int_{0}^{\mathop{\rm diam}\nolimits(Y)}\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s){\mathrm{d}}s}$. Note that $\sinh^{N-1}(\sqrt{\frac{-K}{N-1}}s)\leq\frac{e^{\sqrt{(N-1)|K|}s}}{2^{N-1}}$, then
$\mathfrak{m}\big(B(y_{0},\mathop{\rm diam}\nolimits(Y)+i+1)\big)\leq c_{1}\int_{0}^{\mathop{\rm diam}\nolimits(Y)+i+1}\frac{e^{\sqrt{(N-1)|K|}s}}{2^{N-1}}\,{\mathrm{d}}s\leq c_{2}e^{\sqrt{(N-1)|K|}i}$ (2.8)
where $c_{2}=c_{1}\frac{e^{\sqrt{(N-1)|K|}(\mathop{\rm diam}\nolimits(Y)+1)}}{2^{N-1}\sqrt{{(N-1)}|K|}}$.
For $t<\frac{D+1}{\sqrt{(N-1)|K|}}$, it holds $\sqrt{(N-1)|K|}-\frac{2(D+1)}{t}<-\frac{D+1}{t}$, so by (2.7) and (2.8),
$\int_{X}e^{\frac{2\psi_{*}}{t}}\,{\mathrm{d}}\mathfrak{m}\leq\sum_{i=0}^{+\infty}e^{\frac{2\sup_{Y}\psi}{t}}c_{2}e^{\frac{-2(D+1)i}{t}}e^{\sqrt{(N-1)|K|}i}\leq\sum_{i=0}^{+\infty}e^{\frac{2\sup_{Y}\psi}{t}}c_{2}e^{\frac{-(D+1)i}{t}}<+\infty.$ (2.9)
Similarly, we can prove $\int_{X}v^{2}e^{\frac{2\psi_{*}}{t}}\,{\mathrm{d}}\mathfrak{m}<+\infty$ and we prove the claim.
By the regularization of heat flow [3, Theorem 6.5], we have $H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}}),H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\cap L^{\infty}(\mathfrak{m})$. Moreover, by the heat kernel estimate (2.2),
$H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\geq\int_{Y}e^{\frac{\psi(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)>0.$ (2.10)
So $\frac{1}{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap L^{\infty}(\mathfrak{m})$ and by chain rule (see [14, Theorem 4.3.3]), we know the function $x\mapsto\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$. ∎
Lemma 2.3.
For any $v\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$ and $\psi\in\mathop{\rm Lip}\nolimits(Y,{\mathrm{d}})$, we have
[…]
For abbreviation, we can write
$\nabla{\mathrm{K}}_{t}[\psi_{*}]=-\mu^{t}[\psi_{*}],\quad\langle D^{2}{\mathrm{K}}_{t}[\psi_{*}]v,v\rangle=-\frac{1}{t}\int_{S}{\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(v)\,{\mathrm{d}}\rho(x).$
Proof.
Let $|s|\leq 1$. By direct computation,
$\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Phi_{t}[\psi_{*}+sv]=-\frac{\int_{X}v(y)e^{\frac{\psi_{*}(y)+sv(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)+sv(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}.$
Similar to Lemma 2.2, we can see that $\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Phi_{t}[\psi_{*}+sv]\right|$ is uniformly bounded in $s$.
By differentiating under the integral defining ${\mathrm{K}}_{t}[\psi_{*}]$ and letting $s=0$, we have
[…]
Similarly, we can prove
$\frac{{\mathrm{d}}^{2}}{{\mathrm{d}}s^{2}}\Phi_{t}[\psi_{*}+sv]$
$=$ $-\frac{1}{t}\left(\frac{\int_{X}v^{2}(y)e^{\frac{\psi_{*}(y)+sv(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)+sv(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}-\left|\frac{\int_{X}v(y)e^{\frac{\psi_{*}(y)+sv(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)+sv(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}\right|^{2}\right)$
and
[…]
2.3 Gradient estimate
To leverage the Hessian formula in Lemma 2.3, we need to bound the variance term from below; this amounts to proving strong concavity of the regularised Kantorovich functional. We achieve this by establishing a gradient estimate for the marginal density $x\mapsto\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)$.
The following auxiliary lemma will be used throughout this section. We refer to [13, 14] for a comprehensive introduction to $L^{2}$-normed tangent module $L^{2}(TX)$. By [19, Proposition 2.9] and [13, Theorem 1.4.11], elements of $L^{2}(TX)$ admit local-coordinate representations. Readers unfamiliar with non-smooth calculus may safely interpret the proof in the language of Riemannian geometry.
Lemma 2.4.
Let $f\in L^{2}(\mathfrak{m})\cap L^{\infty}(\mathfrak{m})$, and $g\in L^{0}(X\times X)$ be with $g(\cdot,y),g(x,\cdot)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$ for $\mathfrak{m}$-a.e. $x,y\in X$. Then $\int_{X}f(y)g(x,y)\,{\mathrm{d}}\mathfrak{m}(y)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$ and for any $\varphi\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$, we have
$\int\langle\nabla_{x}\int_{X}f(y)g(x,y)\,{\mathrm{d}}\mathfrak{m}(y),\varphi(x)\rangle\,{\mathrm{d}}\mathfrak{m}(x)=\int_{X}f(y)\int\langle\nabla g(\cdot,y),\varphi\rangle\,{\mathrm{d}}\mathfrak{m}\,{\mathrm{d}}\mathfrak{m}(y)$ (2.11)
and
$\left|\nabla_{x}\int_{X}f(y)g(x,y)\,{\mathrm{d}}\mathfrak{m}(y)\right|\leq\int_{X}|f(y)||\nabla_{x}g(x,y)|\,{\mathrm{d}}\mathfrak{m}(y).$ (2.12)
In particular, we can write
$\int_{X}f(y)\nabla_{x}g(x,y)\,{\mathrm{d}}\mathfrak{m}(y)=\nabla_{x}\int_{X}f(y)g(x,y)\,{\mathrm{d}}\mathfrak{m}(y)\in L^{2}(TX).$ (2.13)
Proof.
For $\varphi\in{\rm TestF}:=\big\{\phi\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})\cap L^{\infty}(\mathfrak{m})\cap{\mathrm{D}}(\Delta):\Delta\phi\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})\cap L^{\infty}(\mathfrak{m})\}$, by integration by parts and Fubini theorem, we get
$\int_{X}\Delta\varphi(x)\left(\int_{X}f(y)g(x,y)\,{\mathrm{d}}\mathfrak{m}(y)\right)\,{\mathrm{d}}\mathfrak{m}(x)$ (2.14)
$=$ $\int_{X}f(y)\int_{X}g(x,y)\Delta\varphi(x)\,{\mathrm{d}}\mathfrak{m}(x)\,{\mathrm{d}}\mathfrak{m}(y)$
$=$ $-\int_{X}f(y)\int_{S}\langle\nabla_{x}g(x,y),\nabla\varphi(x)\rangle\,{\mathrm{d}}\mathfrak{m}(x)\,{\mathrm{d}}\mathfrak{m}(y).$
By density of $\rm TestF$} (cf. [13, §3.2]) and the Riesz representation theorem for Hilbert module (cf.[14, Theorem 3.2.14, Example 3.2.15 ]), we prove the lemma. ∎
Lemma 2.5.
For any $\psi\in\mathop{\rm Lip}\nolimits(Y,{\mathrm{d}})$, it holds
$|\nabla_{x}\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)|^{2}\leq{\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(v){\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)$ (2.15)
for $\mathfrak{m}$-a.e. $x\in S$, where
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big):=\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\left(\big|\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot))\big|^{2}\right)$
and
$\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)=\frac{\int_{X}e^{\frac{\psi_{*}(z)}{t}}\nabla_{x}p_{\frac{t}{2}}(x,z)\,{\mathrm{d}}\mathfrak{m}(z)}{\int_{X}e^{\frac{\psi_{*}(z)}{t}}p_{\frac{t}{2}}(x,z)\,{\mathrm{d}}\mathfrak{m}(z)}.$
Proof.
First of all, by [22], $p_{s}(\cdot,y)\in W^{1,2}(X,{\mathrm{d}},\mathfrak{m})$ for $\mathfrak{m}$-a.e. $y\in X$ so all the formulas above are well-defined. By Lemma 2.2, Lemma 2.4 and the chain rule for $L^{2}$-normed modules (cf. [14, Theorem 4.3.3]), we have
$\nabla_{x}\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)=\frac{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\nabla H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})-H_{\frac{t}{2}}(ve^{\frac{\psi_{*}}{t}})\nabla H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})}{\big(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\big)^{2}}(x)$
$=$ $\int_{X}v(y)\nabla_{x}\log p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\int_{X}\nabla_{x}\log p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)$
$=$ $\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\left(\nabla_{x}\log p_{\frac{t}{2}}(x,y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\right)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y),$
where in the last equality we use the identity
$\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)$
$=$ $\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\int_{X}\big(v(y)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)\big)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)=0.$
By Lemma 2.4 and Hölder inequality we get (2.15). ∎
Lemma 2.6.
The following identity holds:
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)=\Delta_{x}\log\left(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\right)-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\Delta_{x}\log p_{\frac{t}{2}}(x,\cdot)\big).$ (2.16)
Proof.
By chain rule of Laplacian (cf. [14, Theorem 5.2.3]), it holds
$\Delta\log u=\frac{\Delta u}{u}-|\nabla\log u|^{2},$ (2.17)
for $u=p_{\frac{t}{2}}(x,y)$ and $H_{\frac{t}{2}}(e^{\frac{b}{t}})$. By Lemma  2.4, we have
${\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)$ (2.18)
$=$ $\int_{X}|\nabla_{x}\log p_{\frac{t}{2}}(x,y)|^{2}\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)-\left|\int_{X}\nabla_{x}\log p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)\right|^{2}$
[…]
$-\left|\int_{X}\nabla_{x}\log p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)\right|^{2}$
$=$ $\frac{\Delta_{x}H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)}{H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)}-|\nabla\log H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)|^{2}-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\Delta_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)$
[…]
which is the thesis. ∎
Lemma 2.7.
For $t>0$ small enough, it holds
$J_{t}(x):=\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(|\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)|^{2}\big)\leq\frac{C_{1}(K,N,\Lambda_{\psi})}{t^{2}}.$ (2.19)
Proof.
Integrating the following Li–Yau type estimate (cf. [21, Theorem 1.2])
$|\nabla_{x}\log p_{\frac{t}{2}}(x,y)|^{2}\leq e^{-\frac{Kt}{3}}\frac{\Delta_{x}p_{\frac{t}{2}}(x,y)}{p_{\frac{t}{2}}(x,y)}+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}$ (2.20)
with respect to $\mu_{x}^{t}[\psi_{*}]$, we obtain
$J_{t}(x)\leq e^{-\frac{Kt}{3}}\int_{X}\frac{\Delta_{x}p_{\frac{t}{2}}(x,y)}{p_{\frac{t}{2}}(x,y)}\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}]+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}.$ (2.21)
By symmetry of the heat kernel, we have
$\int_{X}\Delta_{x}p_{\frac{t}{2}}(x,y)e^{\frac{\psi_{*}(y)}{t}}\,{\mathrm{d}}\mathfrak{m}(y)=\int_{X}\Delta_{y}p_{\frac{t}{2}}(x,y)e^{\frac{\psi_{*}(y)}{t}}\,{\mathrm{d}}\mathfrak{m}(y).$
Then by integration by parts formula,
$\int_{X}\frac{\Delta_{x}p_{\frac{t}{2}}(x,y)}{p_{\frac{t}{2}}(x,y)}\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)=$ $\frac{\int_{X}\Delta_{x}p_{\frac{t}{2}}(x,y)e^{\frac{\psi_{*}(y)}{t}}\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}$ (2.22)
$=$ $\frac{\int_{X}\Delta_{y}p_{\frac{t}{2}}(x,y)e^{\frac{\psi_{*}(y)}{t}}\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}$
$=$ $\frac{-\int_{X}\langle\nabla_{y}p_{\frac{t}{2}}(x,y),\nabla_{y}e^{\frac{\psi_{*}(y)}{t}}\rangle\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}$
$\leq$ $\int_{X}|\nabla_{y}\log p_{\frac{t}{2}}(x,y)|\left|\frac{\nabla_{y}\bar{\psi}(y)-\Lambda_{\psi}\nabla_{y}{\mathrm{d}}(y,Y)}{t}\right|\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)$
$\leq$ $\frac{2\Lambda_{\psi}}{t}\left(\underbrace{\int_{X}|\nabla_{y}\log p_{\frac{t}{2}}(x,y)|^{2}{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)}_{:=\tilde{J}_{t}(x)}\right)^{\frac{1}{2}}.$
Similarly, we can prove
$\bar{J}_{t}(x)\leq e^{-\frac{Kt}{3}}\frac{2\Lambda_{\psi}}{t}\bar{J}_{t}^{\frac{1}{2}}(x)+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}.$ (2.23)
Since $1-e^{-\frac{Kt}{3}}=\frac{K}{3}t+o(t)$, we have $\bar{J}_{t}(x)\lesssim\frac{1}{t^{2}}$. Then (2.21) and (2.22) implies
$J_{t}(x)\leq e^{-\frac{Kt}{3}}\frac{2\Lambda_{\psi}}{t}\bar{J}_{t}^{\frac{1}{2}}(x)+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}\leq\frac{C_{1}(K,N,\Lambda_{\psi})}{t^{2}}.$ (2.24)
Lemma 2.8.
Let $x_{0}\in S$, $0<r_{0}\leq 1$ and $B=B(x_{0},r_{0})\subseteq S$. For $t>0$ small enough, it holds
$\int_{B}{\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\rho(x)\leq\frac{C_{2}\mathfrak{m}(B(x_{0},2r_{0}))}{r_{0}t},$ (2.25)
where $C_{2}$ depends on $K,N,\Lambda_{\psi},a_{2}$.
Proof.
[…]
$\int_{B}{\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\rho(x)$ (2.26)
$\leq$ $a_{2}\int_{X}\eta(x){\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}\big(\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\mathfrak{m}(x)$
$=$ $a_{2}\left(\int_{X}\eta(x)\Delta_{x}\log\left(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})(x)\right)\,{\mathrm{d}}\mathfrak{m}(x)-\int_{X}\eta(x)\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\big(\Delta_{x}\log p_{\frac{t}{2}}(x,\cdot)\big)\,{\mathrm{d}}\mathfrak{m}(x)\right)$
= $a_{2}(I_{1}+I_{2}).$
For $I_{1}$, by integration by parts formula and Hölder inequality,
$I_{1}$ $=-\int_{X}\langle\nabla\eta,\nabla\log\big(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\big)\rangle\,{\mathrm{d}}\mathfrak{m}$ (2.27)
$\leq\left(\int_{B(x_{0},2r_{0})\setminus B}|\nabla\eta|^{2}\,{\mathrm{d}}\mathfrak{m}\right)^{\frac{1}{2}}\left(\int_{B(x_{0},2r_{0})\setminus B}\left|\nabla\log\big(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\big)\right|^{2}\,{\mathrm{d}}\mathfrak{m}\right)^{\frac{1}{2}}.$
By Lemma 2.4, it holds
$\left|\nabla\log\big(H_{\frac{t}{2}}(e^{\frac{\psi_{*}}{t}})\big)\right|(x)\leq\int_{X}|\nabla_{x}\log p_{\frac{t}{2}}(x,y)|\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y).$ (2.28)
Combining with Lemma 2.7, we get
$I_{1}\leq\frac{10\sqrt{C_{1}}\mathfrak{m}(B(x_{0},2r_{0}))}{r_{0}t}.$ (2.29)
For $I_{2}$, by (2.17) and the Li–Yau type estimate (2.20), we have
$-\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(\Delta_{x}\log p_{\frac{t}{2}}(x,\cdot))$ $=\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}\left(|\nabla_{x}\log p_{\frac{t}{2}}(x,\cdot)|^{2}-\frac{\Delta_{x}p_{\frac{t}{2}}(x,\cdot)}{p_{\frac{t}{2}}(x,\cdot)}\right)$ (2.30)
$\leq(e^{-\frac{Kt}{3}}-1)\int_{X}\frac{\Delta_{x}p_{\frac{t}{2}}(x,y)}{p_{\frac{t}{2}}(x,y)}\,{\mathrm{d}}\mu_{x}^{t}[\psi_{*}](y)+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}.$
By (2.22) and (2.23), for $t$ small enough, there is $c_{1}=c_{1}(K,N,\Lambda_{\psi})$ so that
$\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(\Delta_{x}\log p_{\frac{t}{2}}(x,y))\leq\left|e^{-\frac{Kt}{3}}-1\right|\frac{2\Lambda_{\psi}}{t}\bar{J}_{t}^{\frac{1}{2}}(x)+\frac{NK}{3}\frac{e^{-\frac{2Kt}{3}}}{1-e^{-\frac{Kt}{3}}}\leq\frac{c_{1}}{t}.$
So $I_{2}\leq\frac{c_{1}\mathfrak{m}(B(x_{0},2r_{0}))}{t}$. Combining with (2.26), (2.29), we obtain (2.25) ∎
Proposition 2.9.
It holds that
$\int_{B}|\nabla_{x}\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)|\,{\mathrm{d}}\rho(x)\leq\frac{\sqrt{C_{2}\mathfrak{m}(B(x_{0},2r_{0}))}}{\sqrt{r_{0}t}}\left(\int_{B}{\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(v)\,{\mathrm{d}}\rho(x)\right)^{\frac{1}{2}}.$ (2.31)
Proof.
It follows from Lemma 2.5, Lemma 2.8, and Hölder inequality. ∎
2.4 Global concavity estimate
We improve the estimate (2.31) in Proposition2.9 to a global one. The proof is similar to [30, Lemma 3.3], which involves two key lemmas in Appendix A.
Proposition 2.10.
It holds that
$\int_{S}\big|\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)-\mathbb{E}_{\mu^{t}[\psi_{*}]}(v))\big|\,{\mathrm{d}}\rho(x)\leq\frac{\kappa}{\sqrt{t}}\left(\int_{S}{\rm{Var}}_{\mu_{x}^{t}[\psi_{*}]}(v)\,{\mathrm{d}}\rho(x)\right)^{\frac{1}{2}},$ (2.32)
where $\mu^{t}[\psi_{*}]$ is defined as in (2.6) and $\kappa$ depends on $K,N,\Lambda_{\psi},a_{1},a_{2},S$.
Proof.
[…]
$\int_{S}\big|\mathbb{E}_{\mu_{x}^{t}[\psi_{*}]}(v)-\mathbb{E}_{\mu^{t}[\psi_{*}]}(v))\big|\,{\mathrm{d}}\rho(x)$ (2.33)
[…]
which is the thesis. ∎
2.5 Stability estimate for Kantorovich functionals
Proposition 2.11.
Let $\varphi,\psi\in\mathop{\rm Lip}\nolimits(Y,{\mathrm{d}})$ and $t>0$. Then
$\int_{S}\left|\Big(\Phi_{t}[\varphi_{*}](x)-\mathrm{K}_{t}[\varphi_{*}]\Big)-\Big(\Phi_{t}(\psi_{*})(x)-\mathrm{K}_{t}[\psi_{*}]\Big)\right|\,{\mathrm{d}}\rho(x)$ (2.34)
$\leq$ $C_{5}\left|\mathbb{E}_{\mu^{t}[\varphi_{*}]-\mu^{t}[\psi_{*}]}(\varphi_{*}-\psi_{*})\right|^{\frac{1}{2}},$
where $C_{5}$ depends on $K,N,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),S,\mathop{\rm Lip}\nolimits(\varphi),\mathop{\rm Lip}\nolimits(\psi)$.
Proof.
Denote $v=\varphi_{*}-\psi_{*}$, and $\phi_{s}=\psi_{*}+sv$ for $0\leq s\leq 1$. By Lemma 2.3,
$\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Big({\Phi}_{t}[\phi_{s}]-\mathrm{K}_{t}[\phi_{s}]\Big)=$ $-\left(\mathbb{E}_{\mu_{x}^{t}[\phi_{s}]}(v)-\mathbb{E}_{\mu^{t}[\phi_{s}]}(v)\right).$ (2.35)
By Proposition 2.10, we have
$\int_{S}\left|\Big(\Phi_{t}[\varphi_{*}](x)-\mathrm{K}_{t}[\varphi_{*}]\Big)-\Big(\Phi_{t}[\psi_{*}](x)-\mathrm{K}_{t}[\psi_{*}]\Big)\right|\,{\mathrm{d}}\rho(x)$
$\leq$ $\int_{0}^{1}\int_{S}\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Big({\Phi}_{t}[\phi_{s}]-\mathrm{K}_{t}[\phi_{s}]\Big)\right|\,{\mathrm{d}}\rho(x)\,{\mathrm{d}}s=\int_{0}^{1}\int_{S}\left|\mathbb{E}_{\mu_{x}^{t}[\phi_{s}]}(v)-\mathbb{E}_{\mu^{t}[\phi_{s}]}(v)\right|\,{\mathrm{d}}\rho(x)\,{\mathrm{d}}s$
$\overset{*}{\leq}$ $\kappa\left(-\int_{0}^{1}\langle D^{2}{\mathrm{K}}_{t}[\phi_{s}]v,v\rangle\,{\mathrm{d}}s\right)^{\frac{1}{2}}$
$=$ $\kappa\left|\langle\nabla{\mathrm{K}}_{t}[\varphi_{*}]-\nabla{\mathrm{K}}_{t}[\psi_{*}],\varphi_{*}-\psi_{*}\rangle\right|^{\frac{1}{2}}=\kappa\left|\mathbb{E}_{\mu^{t}[\varphi_{*}]-\mu^{t}[\psi_{*}]}(\varphi_{*}-\psi_{*})\right|^{\frac{1}{2}},$
where ()(*) follows from Proposition 2.10, Lemma 2.3 and Hölder inequality, while the last two equalities follow from Lemma 2.3. This completes the proof. ∎
2.6 Passing to the limit
In the following lemma we pass to the limit of $t$ to recover the Kantorovich potentials. We remark that, for $\psi\in C_{b}(X)$, this asymptotic formula has been proved by Gigli–Tamanini–Trevisan [17, Proposition 5.2] in ${\rm RCD}(K,\infty)$ spaces.
Lemma 2.12.
For any $\psi\in\mathop{\rm Lip}\nolimits(Y,{\mathrm{d}})$, we have
$\lim_{t\rightarrow 0}{\Phi}_{t}[\psi_{*}](x)=\lim_{t\rightarrow 0}-t\log\int_{X}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)=\psi^{c}(x),$ (2.36)
for $\rho$-a.e. $x\in S$, where $\psi^{c}(x)=\inf_{y\in Y}\{c(x,y)-\psi(y)\}$.
Proof.
For $\delta>0$ and $Y_{\delta}:=\{y\in X:{\mathrm{d}}(y,Y)\leq\delta\}$, we denote
$\int_{X}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)=\underbrace{\int_{Y_{\delta}}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}_{I_{1}}+\underbrace{\int_{X\setminus Y_{\delta}}e^{\frac{\psi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}_{I_{2}}.$ (2.37)
Step 1: Recall that $\psi_{*}(y)=\bar{\psi}(y)-\Lambda_{\psi}{\mathrm{d}}(y,Y)$, $\Lambda_{\psi}=D+\mathop{\rm Lip}\nolimits(\psi)+1,D=\mathop{\rm diam}\nolimits(S\cup Y)$. We have
$\frac{I_{2}}{I_{1}}\leq$ $\frac{e^{\frac{\sup_{Y}\psi_{i}-(D+1)\delta}{t}}\int_{X}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{e^{\frac{\inf_{Y}\psi_{i}}{t}}\int_{Y}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}.$ (2.38)
Combining with the stochastic completeness of the heat flow and the heat kernel estimate (2.10), we can find $\delta>0$ large enough such that
$\frac{I_{2}}{I_{1}}\lesssim e^{-\frac{D^{2}}{t}}~~~~\text{as}~~t\to 0.$ (2.39)
Step 2: We claim that
$\inf_{y\in Y_{\delta}\setminus Y}\{c(x,y)-\psi_{*}(y)\}>\inf_{y\in Y}\{c(x,y)-\psi(y)\}=\psi^{c}(x).$ (2.40)
Since $Y$ is compact, there exists $y_{1}\in\overline{Y_{\delta}\setminus Y}$ and $y_{2}\in Y$, such that
$c(x,y_{1})-\psi_{*}(y_{1})=\inf_{y\in Y_{\delta}\setminus Y}\{c(x,y)-\psi_{*}(y)\},~~~{\mathrm{d}}(y_{1},y_{2})={\mathrm{d}}(y_{1},Y)=:s>0.$
Then (2.40) follows from the following estimate
$c(x,y_{1})-\psi_{*}(y_{1})-\big(c(x,y_{2})-\psi(y_{2})\big),$ (2.41)
$\geq$ $\frac{1}{2}\big({\mathrm{d}}(x,y_{2})-{\mathrm{d}}(y_{1},y_{2})\big)^{2}-\frac{1}{2}{\mathrm{d}}^{2}(x,y_{2})-\mathop{\rm Lip}\nolimits(\bar{\psi}){\mathrm{d}}(y_{1},y_{2})+\Lambda_{\psi}{\mathrm{d}}(y_{1},y_{2})$
$\geq$ $\frac{1}{2}s^{2}+(\Lambda_{\psi}-\mathop{\rm Lip}\nolimits(\psi)-D)s=\frac{1}{2}s^{2}+s>0.$
Step 3: It holds that
$\lim_{t\rightarrow 0}-t\log I_{1}=\psi^{c}.$ (2.42)
Then we obtain (2.36) by combing (2.39) and (2.42).
Lower bound:
[…]
Denote $c_{t}(x,y)=-t\log p_{\frac{t}{2}}(x,y)$. By Lemma 2.1, we have $c_{t}(x,y)\rightarrow c(x,y)$ uniformly in $S\times Y_{\delta}$. So for any $\epsilon>0$, there is $t_{0}>0$, such that for any $t<t_{0}$, $c_{t}(x,y)\geq c(x,y)-\epsilon$, and
$-t\log\int_{Y_{\delta}}e^{\frac{\psi_{*}(y)-c_{t}(x,y)}{t}}{\mathrm{d}}\mathfrak{m}(y)\geq$ $-t\log\int_{Y_{\delta}}e^{\frac{\psi_{*}(y)-c(x,y)+\epsilon}{t}}\,{\mathrm{d}}\mathfrak{m}(y)$ (2.44)
[…]
$=$ $-t\left(\log\mathfrak{m}(Y_{\delta})+\frac{-\psi^{c}(x)+\epsilon}{t}\right)$
$=$ $\psi^{c}(x)-t\log\mathfrak{m}(Y_{\delta})-\epsilon.$
Letting $t\rightarrow 0$, then letting $\epsilon\rightarrow 0$, we get (2.43).
Upper bound:
[…]
For any $x\in S$, by compactness of $Y$, there exists $T(x)\in Y$, such that
$\psi(T(x))+\psi^{c}(x)=c(x,T(x)).$
For any $\epsilon>0$, by continuity of $\psi_{*}$ and $c(x,y)$, there is $r<\delta\wedge\epsilon$ such that
$\psi_{*}(y)\geq\psi(T(x))-\epsilon,\quad c(x,y)\leq c(x,T(x))+\epsilon,~~~\forall y\in B(T(x),r).$
Moreover, there exists $t_{0}>0$, such that for any $t<t_{0}$, there holds
$c_{t}(x,y)\leq c(x,y)+\epsilon.$
Then for any $t<t_{0}$, we have
$-t\log\int_{Y_{\delta}}e^{\frac{\psi_{*}(y)-c_{t}(x,y)}{t}}\,{\mathrm{d}}\mathfrak{m}(y)\leq$ $-t\log\int_{B(T(x),r)}e^{\frac{\psi(T(x))-c(x,T(x))-3\epsilon}{t}}\,{\mathrm{d}}\mathfrak{m}(y)$ (2.46)
$=$ $-t\log\int_{B(T(x),r)}e^{\frac{-\psi^{c}(x)-3\epsilon}{t}}\,{\mathrm{d}}\mathfrak{m}(y)$
$=$ $-t\log\mathfrak{m}(B(T(x),r))+\psi^{c}(x)+3\epsilon.$
Letting $t\rightarrow 0$, then $\epsilon\rightarrow 0$, we get (2.45).
Lemma 2.13.
Let $\varphi$ be a Kantorovich potentials from $\nu$ to $\rho$. Then $\mu^{t}[\varphi_{*}]\rightharpoonup\nu$, i.e., for any $v\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$, it holds
$\lim_{t\rightarrow 0}\mathbb{E}_{\mu^{t}[\varphi_{*}]}(v)=\mathbb{E}_{\nu}(v).$ (2.47)
Proof.
We just need to show that
$\lim_{t\rightarrow 0}\mathbb{E}_{\mu_{x}^{t}[\varphi_{*}]}(v)=v(T(x)),\quad\text{for $\rho$-a.e. $x\in S$},$ (2.48)
where $T$ is the unique optimal transport map from $\rho$ to $\nu$. The proof is very similar to that of Lemma 2.12, so we only sketch it.
Step 1: Without loss of generality, we assume that $v\geq 0$. Similar to (2.39), there is $\delta>0$ such that
$\frac{\int_{X\setminus Y_{\delta}}v(y)e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}\lesssim e^{-\frac{D^{2}}{t}}~~\text{as}~~t\to 0.$ (2.49)
and
$\frac{\int_{X\setminus Y_{\delta}}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}\lesssim e^{-\frac{D^{2}}{t}}~~\text{as}~~t\to 0.$ (2.50)
Step 2: By uniqueness of optimal transport $T$ and (2.40), for almost every $x\in S$, there is a unique $T(x)\in Y$ so that
$\varphi^{c}(x)=c(x,T(x))-\varphi_{*}(T(x))=\inf_{y\in Y_{\delta}}\big\{c(x,y)-\varphi_{*}(y)\big\}.$
Then for any $\epsilon>0$, we can find $0<r_{2}<\delta$ such that
$|v(y)-v(T(x)|<\epsilon,~~~~~\forall y\in B\big(T(x),r_{2}\big)$ (2.51)
and
$\varphi_{*}(y)-c(x,y)<-\varphi^{c}-4\delta_{r_{2}},~~~~~\forall y\in Y_{\delta}\setminus B\big(T(x),r_{2}\big)$ (2.52)
for some $\delta_{r_{2}}>0$. By Lemma 2.1, there is $t_{0}>0$ such that
$|c_{t}(x,y)-c(x,y)|<\delta_{r_{2}},~~~\forall(x,y)\in S\times Y~\text{and}~t<t_{0}.$
By the same argument as (2.44), for $t<t_{0}$ we have
${e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)}\leq{e^{\frac{-\varphi^{c}(x)-3\delta_{r_{2}}}{t}}},~~~\forall y\in Y_{\delta}\setminus B\big(T(x),r_{2}\big),$ (2.53)
Furthermore, similar to (2.46), we can find $r_{1}<r_{2}$ so that
${e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)}>{e^{\frac{-\varphi^{c}(x)-2\delta_{r_{2}}}{t}}},~~~\forall y\in B\big(T(x),r_{1}\big)~\text{and}~t<t_{0}.$ (2.54)
It follows from (2.53), (2.54) that
$\frac{\int_{Y_{\delta}\setminus B(T(x),r_{2})}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{B(T(x),r_{2})}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}\to 0~~\text{as}~~t\to 0.$ (2.55)
Combining (2.49), (2.50), (2.51) and (2.55) we obtain
$\lim_{t\rightarrow 0}|\mathbb{E}_{\mu_{x}^{t}[\varphi_{*}]}(v)-v(T(x))|$
$=$ $\lim_{t\rightarrow 0}\left|\frac{\int_{X}\big(v(y)-v(T(x))\big)e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{X}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}\right|$
$=$ $\lim_{t\rightarrow 0}\left|\frac{\int_{B(T(x),r_{2})}\big(v(y)-v(T(x))\big)e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}{\int_{B(T(x),r_{2})}e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)\,{\mathrm{d}}\mathfrak{m}(y)}\right|$
$\leq$ $\epsilon.$
Letting $\epsilon\to 0$ we get (2.48).
Proof of Theorem 1.1.
Let $\varphi$ and $\psi$ be Kantorovich potentials from $\mu$ to $\rho$ and $\nu$ to $\rho$ respectively. By the heat kernel lower bound (2.10) and the stochastic completeness of the heat flow, we can see that $|\Phi_{t}[\psi_{*}]|,|\Phi_{t}[\varphi_{*}]|$ are uniformly bounded for $t\in(0,1]$. So by the dominated convergence theorem and Lemma 2.12 we obtain
$\lim_{t\rightarrow 0}\int_{S}\left|\Big(\Phi_{t}[\varphi_{*}](x)-\mathrm{K}_{t}[\varphi_{*}]\Big)-\Big(\Phi_{t}[\psi_{*}](x)-\mathrm{K}_{t}[\psi_{*}]\Big)\right|\,{\mathrm{d}}\rho(x)$ (2.56)
$=$ $\int_{S}\left|\Big(\varphi^{c}(x)-\mathbb{E}_{\rho}(\varphi^{c})\Big)-\Big(\psi^{c}(x)-\mathbb{E}_{\rho}(\psi^{c})\Big)\right|\,{\mathrm{d}}\rho(x).$
[…]
$\lim_{t\rightarrow 0}\mathbb{E}_{\mu^{t}[\varphi_{*}]-\mu^{t}[\psi_{*}]}(\varphi_{*}-\psi_{*})=\mathbb{E}_{\mu-\nu}(\varphi-\psi).$ (2.57)
Combining Proposition 2.11, (2.56) and (2.57), we obtain
$\int_{S}\left|\Big(\varphi^{c}(x)-\mathbb{E}_{\rho}(\varphi^{c})\Big)-\Big(\psi^{c}(x)-\mathbb{E}_{\rho}(\psi^{c})\Big)\right|\,{\mathrm{d}}\rho(x)\leq\sqrt{C\left|\mathbb{E}_{\mu-\nu}(\varphi-\psi)\right|}$ (2.58)
for some $C>0$. Note that $c(x,y)=\frac{1}{2}{\mathrm{d}}^{2}(x,y)$, thus the Kantorovich potential $\varphi$ and $\psi$ are $\mathop{\rm diam}\nolimits(S\cup Y)$-Lipschitz, so $C$ depends only on $K,N,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),S$.
By assumption, $\phi_{\nu}=\varphi^{c},\phi_{\mu}=\psi^{c}$ satisfies $\mathbb{E}_{\rho}(\phi_{\mu})=\mathbb{E}_{\rho}(\phi_{\nu})=0$. Note also that $\mathop{\rm Lip}\nolimits(\varphi-\psi)\leq 2\mathop{\rm diam}\nolimits(S\cup Y)$, by Kantorovich duality for $W_{1}$, we finally obtain
$\|\phi_{\mu}-\phi_{\nu}\|^{2}_{L^{1}(\rho)}\leq 2C\mathop{\rm diam}\nolimits(S\cup Y)W_{1}(\mu,\nu).$ (2.59)
3 Stability of optimal transport maps
In this section, we prove Theorem 1.3 concerning the quantitative stability of optimal transport maps on Alexandrov spaces. This is achieved by combining Theorem 1.1 and the following estimate.
Theorem 3.1.
[…]
$\int_{S}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2}\,{\mathrm{d}}\rho(x)\leq\bar{C}\left(\int_{S}|\phi_{\mu}(x)-\phi_{\nu}(x)|^{2}\,{\mathrm{d}}\rho(x)\right)^{\frac{1}{3}},$ (3.1)
where $\phi_{\mu}$ and $\phi_{\nu}$ are the Kantorovich potentials from $\rho$ to $\mu$ and $\rho$ to $\nu$ respectively.
We follow a strategy of [26] by lifting integrals from the base space to the unit tangent bundle, evolve the relevant quantities via the geodesic flow developed in [25], and then exploit the one-dimensional convexity of the potentials. For the reader’s convenience, we recall the necessary Alexandrov geometry theory and provide a self-contained proof.
The Riemannian structure on the set of regular points $X_{\mathrm{reg}}$ (cf.[38]) endows the tangent bundle $\mathrm{T}X_{\mathrm{reg}}$ an Euclidean vector bundle structure. On this Euclidean vector bundle, one can define (cf. [25, Section3] ) a canonical Liouville measure $\mathfrak{m}_{\mathrm{L}}$: it is the unique Borel measure on $\mathrm{T}X_{\mathrm{reg}}$ such that for any Borel set $A\subset\mathrm{T}X_{\mathrm{reg}}$, we have
$\mathfrak{m}_{\mathrm{L}}(A)=\int_{X}\mathcal{H}^{n}\big(A\cap\mathrm{T}_{x}X\big)\,{\mathrm{d}}\mathcal{H}^{n}(x).$
We extend $\mathfrak{m}_{\mathrm{L}}$ to a measure on the tangent bundle $\mathrm{T}X$ by setting $\mathfrak{m}_{\mathrm{L}}(\mathrm{T}X\setminus\mathrm{T}X_{\mathrm{reg}})=0$. Moreover, by [6, Theorem 1.4], if $(X,{\mathrm{d}})$ is an Alexandrov space with no boundary, then the geodesic flow preserves the Liouville measure.
Let $\mathrm{S}X=\{v\in\mathrm{T}X:|v|=1\}$ denote the unit tangent bundle (sphere bundle). There is a canonical Liouville measure $\mathfrak{m}_{\mathrm{S}}$ on $\mathrm{S}X$ (cf.[25, Section 3.6]), also called Liouville measure, such that if the geodesic flow defined on ${\rm T}X$ preserves $\mathfrak{m}_{\mathrm{L}}$, then the geodesic flow defined on ${\mathrm{S}}X$ preserves $\mathfrak{m}_{\mathrm{S}}$ as well. For any Borel set $A\subseteq{\mathrm{S}}X$, we have
$\mathfrak{m}_{\mathrm{S}}(A)=\int_{{\mathrm{S}}X}\chi_{A}(x,v)\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)=c_{n}\int_{X}\int_{\Sigma_{x}(X)}\chi_{A}(x,v)\,{\mathrm{d}}\sigma_{x}(v)\,{\mathrm{d}}\mathcal{H}^{n}(x),$ (3.2)
where $\Sigma_{x}(X)$ denotes the space of directions at $x$, $c_{n}=\mathcal{H}^{n-1}({\mathrm{S}}^{n-1})$ and $\sigma_{x}\in\mathcal{P}(\Sigma_{x}(X))$ is the canonical probability measure on the fiber.
Lifting to the unit tangent bundle.
Let $\varphi_{s}:{\mathrm{S}}X\rightarrow{\mathrm{S}}X$ be the geodesic flow at time $s\in[0,1]$ on $\mathrm{S}X$, and write $\varphi_{s}(x,v)=(b_{s}(x,v),t_{s}(x,v))$, where $b_{s}(x,v)\in X$ and $t_{s}(x,v)\in{\Sigma_{b_{s}(x,v)}(X)}$. For $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}X$ such that the curve $0,1]\ni s\mapsto b_{s}(x,v)$ is a locally shortest path, denote by $I_{S}(x,v)$ the set of connected components of $\{s\in[0,1]:b_{s}(x,v)\in S\}$. Since $S$ is an open set, we have
$\{s\in[0,1]:b_{s}(x,v)\in S\}=\bigcup_{i\in I_{S}(x,v)}(\alpha_{i}(x,v),\beta_{i}(x,v)).$
For $x\in S\cap X_{\mathrm{reg}}$, by [8, Theorem 10.8.4], $\Sigma_{x}(X)$ is isometric to $\mathrm{S}^{n-1}$, then there is a universal constant $d_{n}$ such that
$\int_{\Sigma_{x}(X)}\langle\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x),v\rangle^{2}\,{\mathrm{d}}\sigma_{x}(v)$ (3.3)
$=$ $\int_{{\mathrm{S}}^{n-1}}\langle\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x),v\rangle^{2}\,{\mathrm{d}}\sigma_{x}(v)$
$=$ $d_{n}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2}.$
By [25, Section 3.6], $\mathfrak{m}_{\mathrm{S}}$ is preserved by $\varphi_{s}$, so
$\int_{S}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2}\,{\mathrm{d}}\rho(x)\leq a_{2}\int_{S}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2}\,{\mathrm{d}}\mathcal{H}^{n}(x)$ (3.4)
[…]
$=$ $\frac{a_{2}}{c_{n}d_{n}}\int_{{\mathrm{S}}X}\int_{0}^{1}\langle\nabla\phi_{\mu}(b_{s}(x,v))-\nabla\phi_{\nu}(b_{s}(x,v)),t_{s}(x,v)\rangle^{2}\chi_{S}(b_{s}(x,v))\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$
$=$ $\frac{a_{2}}{c_{n}d_{n}}\int_{{\mathrm{S}}X}\sum_{i\in I_{S}(x,v)}\int_{\alpha_{i}(x,v)}^{\beta_{i}(x,v)}\langle\nabla\phi_{\mu}(b_{s}(x,v))-\nabla\phi_{\nu}(b_{s}(x,v)),t_{s}(x,v)\rangle^{2}\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v).$
One-dimensional convexity estimate.
For $(x,v)\in{\mathrm{S}}X$ and $s\in[0,1]$, we denote $u_{\mu}^{(x,v)}(s)=\phi_{\mu}(b_{s}(x,v))$. Then
$\frac{{\mathrm{d}}}{{\mathrm{d}}s}u_{\mu}^{(x,v)}(s)=\langle\nabla\phi_{\mu}(b_{s}(x,v)),t_{s}(x,v)\rangle,\quad\text{for a.e. $s\in[0,1]$},$
and an analogous formula holds for $u_{\nu}^{(x,v)}(s)=\phi_{\nu}(b_{s}(x,v))$.
Denote $S_{1}=\{x\in X:{\mathrm{d}}(x,S)\leq 1\}$. Note that in the last integral of (3.4), only the elements $(x,v)\in{\mathrm{S}}X$ for which $x\in S_{1}$ have a non-vanishing contribution. For $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}X$ with $x\in S_{1}$, the curve $s\mapsto b_{s}(x,v)$ is locally minimizing on $[0,1]$. Thus, applying Ohta’s semiconcavity estimate [37, Lemma 3.2] on sufficiently small subsegments yields local concavity of $u^{(x,v)}_{\mu}-\zeta|s|^{2}$ with a uniform constant $\zeta$. The local-to-global principle for concave functions then leads to the following lemma.
Lemma 3.2 ([37], Lemma 3.2).
There exists $\zeta$, depending on $k$ and $\mathop{\rm diam}\nolimits(S\cup Y)$, such that for $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}{S_{1}}$, the functions $u_{\mu}^{(x,v)}-\zeta|s|^{2}$, $u_{\nu}^{(x,v)}-\zeta|s|^{2}$ are concave on $s\in[0,1]$. Moreover, the modulus of their derivatives (which exist a.e. on $[0,1]$) is bounded above by $\mathop{\rm diam}\nolimits(S\cup Y)+2\zeta$.
Applying Lemma 3.3 below to the functions $\zeta|s|^{2}-u_{\mu}^{(x,v)}$ and $\zeta|s|^{2}-u_{\nu}^{(x,v)}$ on each compact segment $\alpha_{i}(x,v),\beta_{i}(x,v)]$ and using Hölder inequality, we obtain
$\int_{S}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2}\,{\mathrm{d}}\rho(x)$ (3.5)
[…]
$\leq$ $C_{1}\int_{{\mathrm{S}}X}\sum_{i\in I_{S}(x,v)}\left(\int_{\alpha_{i}(x,v)}^{\beta_{i}(x,v)}|\phi_{\mu}(b_{s}(x,v))-\phi_{\nu}(b_{s}(x,v))|^{2}\,{\mathrm{d}}s\right)^{\frac{1}{3}}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$
$\leq$ $C_{1}\int_{{\mathrm{S}}X}\left(\#I_{S}\right)^{\frac{2}{3}}\left(\int_{0}^{1}|\phi_{\mu}(b_{s})-\phi_{\nu}(b_{s})|^{2}\chi_{S}(b_{s})\,{\mathrm{d}}s\right)^{\frac{1}{3}}{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}$
$\leq$ $C_{1}\left(\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{2}{3}}\left(\int_{{\mathrm{S}}X}\int_{0}^{1}|\phi_{\mu}(b_{s})-\phi_{\nu}(b_{s})|^{2}\chi_{S}(b_{s})\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{1}{3}}$
$\leq$ $C_{1}\left(\frac{c_{n}}{a_{1}}\right)^{\frac{1}{3}}\left(\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{2}{3}}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}},$
where $C_{1}$ depends on $k,n,a_{2},\mathop{\rm diam}\nolimits(S\cup Y)$, and the last inequality follows from the invariance of $\mathfrak{m}_{\mathrm{S}}$ under the geodesic flow.
Lemma 3.3 ([11], Lemma 5.1).
Let $I\subseteq\mathbb{R}$ be a compact segment and let $u,v:I\rightarrow\mathbb{R}$ be two convex functions such that $|u^{\prime}|$ and $|v^{\prime}|$ (defined a.e. on $I$) are uniformly bounded on $I$. Then
$\|u^{\prime}-v^{\prime}\|_{L^{2}(I)}^{2}\leq 8(\|u^{\prime}\|_{L^{\infty}(I)}+\|v^{\prime}\|_{L^{\infty}(I)})^{\frac{4}{3}}\|u-v\|_{L^{2}(I)}^{\frac{2}{3}}.$
One-dimensional BV estimate.
We now aim to control $\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}$, the average number of times a geodesic crosses the boundary of $S$. The argument of [26, Proposition 4.3] relies on the existence of $T>0$, so that for $(x,v)\in{\mathrm{S}}X$, the geodesic $s\mapsto b_{s}(x,v)$ is minimizing on $[0,T]$, and hence in particular does not self-intersect. In an Alexandrov space, one cannot in general expect such a uniform injectivity radius bound. We use a localization method to overcome this difficulty.
In the following proposition, we denote by $|{\mathrm{D}}f|([0,1])$ the total variation of a function $f$ on $[0,1]$ and denote by ${\rm{Per}}(S)$ the perimeter of a set $S\subset X$. We refer to [33, 1] for $\rm{BV}$ functions and sets of finite perimeter in metric measure spaces.
Proposition 3.4.
Assume ${\rm{Per}}(S)<+\infty$, then
$\int_{{\mathrm{S}}X}\#I_{S}(x,v)\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)\leq c_{n}\big(\mathcal{H}^{n}(S_{1})+{\rm{Per}}(S)\big).$
Proof.
Let $(x,v)\in{\mathrm{S}}X$ be such that $s\mapsto b_{s}(x,v)$ is well-defined on $[0,1]$, and let $u\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$. Note that $s\mapsto b_{s}(x,v)$ has unit speed, then
$\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}u(b_{s}(x,v))\right|\leq\left|\nabla u\right|(b_{s}(x,v)),~~~\text{a.e. $s\in[0,1]$},$
and
$|{\mathrm{D}}u(b_{s}(x,v))|([0,1])=\int_{0}^{1}\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}u(b_{s}(x,v))\right|\,{\mathrm{d}}s\leq\int_{0}^{1}|\nabla u|(b_{s}(x,v))\,{\mathrm{d}}s.$ (3.6)
Let $(u_{k})_{k\in\mathbb{N}}\subseteq\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$ be such that $u_{k}\rightarrow\chi_{S}$ in $L^{1}(\mathcal{H}^{n})$ and
$\int_{X}|\nabla u_{k}|\,{\mathrm{d}}\mathcal{H}^{n}\rightarrow{\rm{Per}}(S).$ (3.7)
Then
$\int_{{\mathrm{S}}X}\int_{0}^{1}\left|u_{k}(b_{s}(x,v))-\chi_{S}(b_{s}(x,v))\right|\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$ (3.8)
$=$ $\int_{0}^{1}\int_{{\mathrm{S}}X}\left|u_{k}(b_{s}(x,v))-\chi_{S}(b_{s}(x,v))\right|\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)\,{\mathrm{d}}s$
${=}$ $c_{n}\int_{X}|u_{k}(x)-\chi_{S}(x)|\,{\mathrm{d}}\mathcal{H}^{n}(x)\rightarrow 0.$
So up to taking a subsequence, it holds that for $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}X$,
$u_{k}(b_{s}(x,v))\rightarrow\chi_{S}(b_{s}(x,v))\quad\text{in $L^{1}([0,1])$ as $k\to\infty$}.$ (3.9)
By lower semicontinuity of the total variation, we have
$\int_{{\mathrm{S}}X}|{\mathrm{D}}\chi_{S}(b_{s}(x,v))|([0,1])\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$ (3.10)
[…]
where ()(*) follows from (3.6) and the invariance of $\mathfrak{m}_{\mathrm{S}}$ under the geodesic flow.
Finally, notice that
$I_{S}(x,v)\leq 1+\frac{1}{2}|{\mathrm{D}}\chi_{S}(b_{s}(x,v))|([0,1])$ (3.11)
and that only the elements $(x,v)\in{\mathrm{S}}X$ for which $x\in S_{1}$ have a non-vanishing contribution. Combining with (3.10), we obtain
$\int_{{\mathrm{S}}X}\#I_{S}(x,v)\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)\leq$ $c_{n}\mathcal{H}^{n}(S_{1})+\frac{1}{2}c_{n}{\rm{Per}}(S).$ (3.12)
Proof of the theorems.
Proof of Theorem 3.1.
By (3.5) and Proposition 3.4, we obtain
$\int_{S}|\nabla\phi_{\mu}-\nabla\phi_{\nu}|^{2}\,{\mathrm{d}}\rho$ (3.13)
$\leq$ $C_{1}\left(\frac{c_{n}}{a_{1}}\right)^{\frac{1}{3}}\left(\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{2}{3}}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}}$
$\leq$ $\bar{C}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}},$
where $\bar{C}$ depends on $k,n,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),{\rm{Per}}(S),S$. ∎
Proof of Theorem 1.3.
By Perelman’s doubling theorem (or Petrunin’s gluing theorem [39]), we may assume that $(X,{\mathrm{d}})$ has no boundary.
Let $\phi_{\mu}$ and $\phi_{\nu}$ be the Kantorovich potentials from $\rho$ to $\mu$ and $\rho$ to $\nu$ respectively. From [4, 41] we know that $\nabla\phi_{\mu}(x),\nabla\phi_{\nu}(x)\in{\rm{T}}_{x}X$ and $T_{\mu}(x)=\exp_{x}(-\nabla\phi_{\mu}(x))$, $T_{\nu}(x)=\exp_{x}(-\nabla\phi_{\nu}(x))$ are well-defined for almost every $x\in S$. By triangle comparison condition (cf. [40]), it holds
${\mathrm{d}}\big(T_{\mu}(x),T_{\nu}(x)\big)\leq c|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|,$ (3.14)
for some constant $c>0$ which depends on $k,\mathop{\rm diam}\nolimits(S\cup Y)$.
Since $\phi_{\mu},\phi_{\nu}$ are $\mathop{\rm diam}\nolimits(S\cup Y)$-Lipschitz and $\mathbb{E}_{\rho}(\phi_{\mu})=\mathbb{E}_{\rho}(\phi_{\nu})=0$, we have
$\|\phi_{\mu}-\phi_{\nu}\|_{L^{\infty}(\rho)}\leq\|\phi_{\mu}\|_{L^{\infty}(\rho)}+\|\phi_{\nu}\|_{L^{\infty}(\rho)}\leq{\rm{osc}}(\phi_{\mu})+{\rm{osc}}(\phi_{\nu})\leq 2\big(\mathop{\rm diam}\nolimits(S\cup Y)\big)^{2}.$
Recall that a finite dimensional Alexandrov space is also $\rm{RCD}$ (cf. [41, 46]). By Theorem 1.1, we get
$\|\phi_{\mu}-\phi_{\nu}\|^{2}_{L^{2}(\rho)}\leq\|\phi_{\mu}-\phi_{\nu}\|_{L^{\infty}(\rho)}\|\phi_{\mu}-\phi_{\nu}\|_{L^{1}(\rho)}\leq C_{3}W_{1}^{\frac{1}{2}}(\mu,\nu).$ (3.15)
Combining Theorem 3.1, (3.14) and (3.15), we obtain
$\int_{S}{\mathrm{d}}^{2}(T_{\mu}(x),T_{\nu}(x))\,{\mathrm{d}}\rho(x)\leq CW_{1}^{\frac{1}{6}}(\mu,\nu),$ (3.16)
where $C$ depends on $k,n,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),{\rm{Per}}(S),S$. This complete the proof. ∎
Appendix A Poincaré inequality: local to global
Lemma A.1.
[…]
$\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}(x)\leq C_{3}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\rho_{B},$ (A.1)
where $C_{3}$ depends on $K,N,a_{1},a_{2}$.
Proof.
By [18, Chapter 9] and [42, Remark 3.3], $\mathfrak{m}$ satisfies the strong local $(1,1)$-Poincaré inequality:
$\int_{B}|f-\mathbb{E}_{\mathfrak{m}_{B}}(f)|\,{\mathrm{d}}\mathfrak{m}\leq c_{1}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\mathfrak{m},$ (A.2)
[…]
$\int_{B}|f-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho\leq$ $\int_{B}|f-\mathbb{E}_{\mathfrak{m}_{B}}(f)|\,{\mathrm{d}}\rho+\int_{B}|\mathbb{E}_{\mathfrak{m}_{B}}(f)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho$ (A.3)
$\leq$ $2a_{2}\int_{B}|f-\mathbb{E}_{\mathfrak{m}_{B}}(f)|\,{\mathrm{d}}\mathfrak{m}$
$\leq$ $2a_{2}c_{1}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\mathfrak{m}\leq 2\frac{a_{2}}{a_{1}}c_{1}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\rho,$
which is the thesis. ∎
Next we prove an $L^{1}$-variant of the gluing Lemma in [30, Lemma 3.3], following the same Boman-chain decomposition, utilizing the doubling property directly instead of the maximal function estimates.
Definition A.2 (Boman chain condition).
We say that a probability measure $\rho$ on an open set $S$ of a metric space satisfies Boman chain condition with parameters $E,F,G>1$ if there is a covering $\mathcal{F}$ of $S$ by open balls $B\in\mathcal{F}$ such that:
  1. 1.
    For any $x\in S$,
    $\sum_{B\in\mathcal{F}}\chi_{2B}(x)\leq E\chi_{S}(x).$
  2. 2.
    For some fixed ball $B_{0}$ in $\mathcal{F}$, called the central ball, and for every $B\in\mathcal{F}$, there exists a chain $B_{0},B_{1},\ldots,B_{N}=B$ of distinct balls from $\mathcal{F}$ such that
    $B\subset FB_{j},~~~\forall j\in\{0,\ldots,N-1\}.$
  3. 3.
    Consecutive balls of the above chain overlap quantitatively:
    $\rho(B_{j}\cap B_{j+1})\geq G^{-1}\max(\rho(B_{j}),\rho(B_{j+1})),~~~\forall j\in\{0,\ldots,N-1\}.$
Lemma A.3.
For any $f\in L^{1}(\rho)$, it holds that
$\int_{S}|f(x)-\mathbb{E}_{\rho}(f)|\,{\mathrm{d}}\rho(x)\leq C_{4}\sum_{B\in\mathcal{F}}\rho(B)\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}(x),$ (A.4)
where $C_{4}$ depends on $K,N,\mathop{\rm diam}\nolimits(S),S$.
Proof.
Since $S$ is a John domain, and $\rho$ is a doubling measure on $S$, by [26, Proposition 3.7], $\rho$ satisfies the Boman chain condition. Hence there exists a covering $\mathcal{F}$ of $S$ satisfying Definition A.2.
For the central ball $B_{0}$, note that
$\int_{S}|f(x)-\mathbb{E}_{\rho}(f)|\,{\mathrm{d}}\rho(x)\leq$ $\int_{S}|f(x)-\mathbb{E}_{\rho_{B_{0}}}(f)|\,{\mathrm{d}}\rho(x)+\int_{S}|\mathbb{E}_{\rho_{B_{0}}}(f)-\mathbb{E}_{\rho}(f)|\,{\mathrm{d}}\rho(x)$ (A.5)
$\leq$ $2\int_{S}|f(x)-\mathbb{E}_{\rho_{B_{0}}}(f)|\,{\mathrm{d}}\rho(x).$
For $B\in\mathcal{F}$, denote $a_{B}=\int_{B}|f-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho=\rho(B)\int_{B}|f-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}$. Since $\mathcal{F}$ is a covering of $S$, we have
$\int_{S}|f(x)-\mathbb{E}_{\rho_{B_{0}}}(f)|\,{\mathrm{d}}\rho(x)\leq\sum_{B\in\mathcal{F}}\int_{B}|f(x)-\mathbb{E}_{\rho_{B_{0}}}(f)|\,{\mathrm{d}}\rho(x)$ (A.6)
$\leq$ $\sum_{B\in\mathcal{F}}\left(\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho(x)+\int_{B}|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\,{\mathrm{d}}\rho(x)\right)$
$\leq$ $\sum_{B\in\mathcal{F}}\left(a_{B}+\rho(B)|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\right).$
For any $B\in\mathcal{F}$, by Boman chain condition, there exists a chain $B_{0},B_{1},\dots,B_{N}=B$ of distinct balls from $\mathcal{F}$, such that for any $j\in\{0,\ldots,N-1\}$,
$|\mathbb{E}_{\rho_{B_{j}}}(f)-\mathbb{E}_{\rho_{B_{j+1}}}(f)|=$ $\left|\frac{1}{\rho(B_{j}\cap B_{j+1})}\int_{B_{j}\cap B_{j+1}}\left(\mathbb{E}_{\rho_{B_{j}}}(f)-\mathbb{E}_{\rho_{B_{j+1}}}(f)\right)\,{\mathrm{d}}\rho\right|$ (A.7)
$\leq$ $\frac{1}{\rho(B_{j}\cap B_{j+1})}\int_{B_{j}\cap B_{j+1}}\left|\mathbb{E}_{\rho_{B_{j}}}(f)-\mathbb{E}_{\rho_{B_{j+1}}}(f)\right|\,{\mathrm{d}}\rho$
$\leq$ $\frac{1}{\rho(B_{j}\cap B_{j+1})}\left(\int_{B_{j}\cap B_{j+1}}\left|f-\mathbb{E}_{\rho_{B_{j}}}(f)\right|+\left|f-\mathbb{E}_{\rho_{B_{j+1}}}(f)\right|\,{\mathrm{d}}\rho\right)$
$\leq$ $\frac{a_{B_{j}}+a_{B_{j+1}}}{\rho(B_{j}\cap B_{j+1})}\overset{*}{\leq}G\left(\frac{a_{B_{j}}}{\rho(B_{j})}+\frac{a_{B_{j+1}}}{\rho(B_{j+1})}\right),$
where ()(*) follows from the quantitative chain overlap of Boman chain condition $3$.
Thus, we have
$|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\leq\sum_{j=0}^{N-1}|\mathbb{E}_{\rho_{B_{j}}}(f)-\mathbb{E}_{\rho_{B_{j+1}}}(f)|\leq 2G\sum_{j=0}^{N}\frac{a_{B_{j}}}{\rho(B_{j})}\overset{*}{\leq}2G\sum_{B\subset F\bar{B}}\frac{a_{\bar{B}}}{\rho(\bar{B})},$ (A.8)
where $\sum_{B\subset F\bar{B}}$ means that the sum runs over all $\bar{B}\in\mathcal{F}$ satisfying $B\subset F\bar{B}$, and ()(*) follows from the Boman chain condition $2$. Then by Fubini–Tonelli theorem,
$\sum_{B\in\mathcal{F}}\rho(B)|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\leq$ $2G\sum_{B\in\mathcal{F}}\rho(B)\sum_{B\subset F\bar{B}}\frac{a_{\bar{B}}}{\rho(\bar{B})}\leq 2G\sum_{\bar{B}\in\mathcal{F}}\frac{a_{\bar{B}}}{\rho(\bar{B})}\sum_{B\subset F\bar{B}}\rho(B).$ (A.9)
By Boman chain condition $1$ and the doubling property of $\rho$, we have
$\sum_{B\subset F\bar{B}}\rho(B)\leq E\rho(F\bar{B})\leq E\beta^{2}F^{\frac{\log\beta}{\log 2}}\rho(\bar{B}),$ (A.10)
where $\beta=\beta(K,N,\mathop{\rm diam}\nolimits(S))$ is the doubling constant.
Combining (A.5), (A.6), (A.9) and (A.10), we obtain
$\int_{S}|f(x)-\mathbb{E}_{\rho}(f)|\,{\mathrm{d}}\rho(x)\leq$ $2\sum_{B\in\mathcal{F}}\left(a_{B}+\rho(B)|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\right)$ (A.11)
$\leq$ $2(1+2\beta^{2}EF^{\frac{\log\beta}{\log 2}}G)\sum_{B\in\mathcal{F}}\rho(B)\int_{B}|f-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B},$
which is the thesis. ∎